OpenStudy (anonymous):

Someone help mewith this quadratic equation pls: Your classmate solved as follows: (x+5)(x-1)=2 x+5=2 x-1=2 x=2-5 x=1+2 x=-3 x=3 Your teacher said this was wrong because the values didnt check. Explain what is wrong with your classmate's solution.

OpenStudy (perl):

more information needed

OpenStudy (anonymous):

It needs a lot more info.

OpenStudy (dan815):

^mmhmm how did he solve it

OpenStudy (anonymous):

ya it really does

OpenStudy (anonymous):

lol sorry there ya go i hit post button too early

OpenStudy (perl):

the problem is , that step is not valid. only if (x-a)(x-b) = 0, can you say that (x-a) =0 or (x-b) = 0. it is false that (x-a) (x-b) = c , that (x-a) = c , (x-b) = c

OpenStudy (perl):

it is false generally * ( it may be true in some fluke example)

OpenStudy (perl):

Or to say it another way, it is false (generally) that the equation (x-a)(x-b)= c implies (x-a) = c or x-b = c It is only true in the case when c = 0

OpenStudy (perl):

so you cannot go from the step ( x+5 )( x -1 ) = 2 to the next step (x+5) =2 , x-1=2. this is not valid . thats why when we solve quadratic equations we set them equal to zero, and then factor

OpenStudy (perl):

the zero product property lets you go from (x+5)(x-2) = 0 to the step (x+5) = 0 or (x-2)=0

OpenStudy (anonymous):

Thank you so much